OpenStudy (anonymous):

Factor the polynomial 3x^2 + x - 10?

OpenStudy (anonymous):

if you do to mathway.com im pretty sure that website could anwser that!

OpenStudy (anonymous):

*go

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\] \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[(x-x_1)(x-x_2)=0\]

OpenStudy (anonymous):

@katiek3455654 it did,thanks (:

OpenStudy (anonymous):

it seriously has helped me out a bunch! np!